# Phantom Menace: Exploring and Enhancing the Robustness of VLA Models against Physical Sensor Attacks

### v1

@@ Introduction
Answer 3 The adversarial-training-based defense indeed enhances the VLA model’s robustness.
As shown in Table 3, adversarial training enhances VLA robustness against physical sensor attacks while preserving clean-data performance. Compared to Table 2, the VLA models experience an average performance decline of approximately 3% on clean datasets. However, for moderate intensity sensor attacks, the model’s performance improves across the board, particularly for OpenVLA, which achieves a maximum performance increase of around 60%.
@@ Conclusion
This paper investigates the robustness of Visual-Language-Action (VLA) models against physical sensor attacks, which is crucial to ensuring their secure deployment in the real world. To achieve efficient and large-scale evaluation, we construct a “Real-Sim-Real” framework that automatically simulates physics-based sensor attack vectors, and validates them on real robotic systems. We also propose an adversarial training defense to mitigate these attacks. We show that existing VLA models are highly vulnerable to physical sensor attacks. Such attacks can severely degrade model performance, resulting in erroneous or hazardous behaviors. Consequently, this vulnerability poses a direct and significant security threat to real-world applications.

### v2

@@ Introduction
Answer 3: The adversarial-training-based defense indeed enhances the VLAs’ robustness.
As shown in Table 3, adversarial training enhances VLA robustness against physical sensor attacks while preserving clean-data performance. Compared to Table 2, the VLA models experience an average performance decline of approximately 3% on clean datasets. However, for moderate-intensity sensor attacks, the model’s performance improves across the board, particularly for OpenVLA, which achieves a maximum performance increase of around 60%.
@@ Conclusion
This paper investigates the robustness of Visual-Language-Action (VLA) models against physical sensor attacks, which is crucial to ensuring their secure deployment in the real world. To achieve efficient and large-scale evaluation, we construct a “Real-Sim-Real” framework that automatically simulates physics-based sensor attack vectors and validates them on real robotic systems. We also propose an adversarial training defense to mitigate these attacks. We show that existing VLA models are highly vulnerable to physical sensor attacks. Such attacks can severely degrade model performance, resulting in erroneous or hazardous behaviors. Consequently, this vulnerability poses a direct and significant security threat to real-world applications.